OpenStudy (anonymous):

The method used to divide complex numbers is similar to the method used to divide radical expressions. Explain why. Give an example.

OpenStudy (anonymous):

Hello - what do you know about the similarities between complex numbers and radical expressions?

OpenStudy (anonymous):

that they are treated the same when solving

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Can you give me an example of a complex number?

OpenStudy (anonymous):

Not doing anything with it, just one unto itself.

OpenStudy (anonymous):

\[\sqrt{-2} = 2i\]

OpenStudy (anonymous):

or i guess \[i \sqrt{2}\]

OpenStudy (anonymous):

That's the imaginary part of a complex number, and yeah, your second one is right. \[ 0+ i\sqrt 2\] What do you notice about the "i"? Might it be a radical? ^_^

OpenStudy (anonymous):

So given the complex number 12+6i, if you divide it by 2i \[\frac{12+6i}{2i}=?\]

OpenStudy (anonymous):

You get \[ \frac{12}{2i}+\frac{6i}{2i} = ?\] What does the second term turn into?

OpenStudy (anonymous):

\[\frac{6}{2} \cdot \cancel{\frac{i}{i}}^1 = ?\]

OpenStudy (anonymous):

For the second expression, the i's cancel out nd leave you with real number. htat's boring though! Looking at the first term \[\frac{12}{2} \cdot \frac{1}{i} = \frac{12}{2} \cdot \frac{1}{\sqrt{-1}}\] And since you have the "i" on the bottom, you have a radical on the bottom - can you rationalize a term with a radical in the denominator?